# LayerSync: Self-aligning Intermediate Layers

### v1

@@ 1 Introduction
Denoising generative models, such as diffusion  (Ho et al., 2020; Song et al., 2020; Song & Ermon, 2019) and flow matching models (Lipman et al., 2023), have demonstrated remarkable success in modeling complex data distributions, achieving state-of-the-art performance across a range of generative tasks. However, this success comes at a significant computation cost. Thus, a new promising line of research has emerged to improve the training efficiency of these models by improving the models’ intermediate representations  (Yu et al., 2024; Wang et al., 2025; Wang & He, 2025). It has been shown that the quality of a diffusion model’s intermediate representations is intrinsically linked to its generative performance. As a result, explicitly guiding these representations accelerates training and improves generation quality  (Yu et al., 2024).
@@
A recent step in this direction is the Dispersive Loss  (Wang & He, 2025), a self-contained regularizer that encourages internal representations to spread out in the feature space, analogous to the repulsive force in contrastive learning  (Oord et al., 2018). Although this approach demonstrates the potential for internal regularization, a substantial performance gap remains compared to methods that leverage external representations. In this paper, we propose a self-contained method with a more directed learning signal to reduce this gap.
Our work is motivated by two key observations: First, while diffusion models learn powerful representations, their quality is highly heterogeneous across the model’s depth. As demonstrated by previous works (Mukhopadhyay et al., 2024; Ghadiyaram, 2025; Stracke et al., 2025) certain intermediate layers capture more semantically rich and useful information than others. Second, when models incorporate knowledge through external guidance using DINOv2 for instance, regularizing early layers seemed more effective than regularizing the deeper ones  (Yu et al., 2024; Wang et al., 2025). Upon these two observations a clear opportunity presents itself: can the model’s own strongest layers act as an intrinsic guidance to improve its weaker ones through self-alignment?
To this end, we propose LayerSync, a simple yet powerful regularization framework that aligns a model’s own intermediate layers. LayerSync is a parameter-free, plug-and-play solution that operates without any external models or data, making it a truly self-contained method. LayerSync introduces negligible computational overhead, yet its effectiveness is substantial. Our experiments show that LayerSync consistently outperforms prior self-contained methods across all tested configurations. For image generation, LayerSync accelerates training on ImageNet 256×256 (Deng et al., 2009) by more than 8.75×. This leads to a new state-of-the-art in purely self-supervised image generation on ImageNet, demonstrating the strength of our self-alignment objective and substantially narrowing the gap between self-contained approaches and those relying on external guidance. Furthermore, due to its self-contained nature, LayerSync seamlessly generalizes to other modalities. Our experiment shows that for audio generation LayerSync leads to $21\%$ improvement in FAD-10K on MTG-Jamendo (Bogdanov et al., 2019), to $7.7\%$ improvement in FID for human motion generation on HumanML3D and $54.7\%$ in FVD for video generation on CLEVRER (see Appendix A). To the best of our knowledge, it is the first time that a self-contained method proves to accelerate diffusion models training seamlessly across different domains.
Additionally, An analysis of the internal features confirms that LayerSync strengthens the model’s representations, leading to a $32.4\%$ improvement in classification and a $63.3\%$ improvement in semantic segmentation.
@@
  • We demonstrate the domain-agnostic versatility of LayerSync by successfully applying it to image, audio, human motion and video generation.
@@ 2 Related Work
Representation learning with diffusion models. Denoising Generative Models including both diffusion  (Ho et al., 2020; Song et al., 2020; Song & Ermon, 2019) and flow matching models (Lipman et al., 2023), trained as multi-level denoising autoencoders (Vincent et al., 2008), naturally give rise to discriminative representations. A line of work has specifically evaluated the quality of these representations, showing that diffusion features can be effectively used across a variety of tasks  (Mukhopadhyay et al., 2024; Ghadiyaram, 2025; Stracke et al., 2025) and, in some cases, achieve performance comparable to self-supervised representation learning methods  (Stracke et al., 2025). However, the quality of the representations varies across model layers, with the final layers, just before the model begins decoding, consistently containing more semantically rich features  Ghadiyaram (2025); Xiang et al. (2023), regardless of whether the architecture is a U-Net  (Ronneberger et al., 2015) or a Transformer  (Vaswani et al., 2017). Our work is directly built upon those insights. We demonstrate that the semantically rich representations in the intermediate layers can be leveraged as a guidance signal to enhance the quality of earlier-layer representations.
@@
Representation regularization for improving diffusion models. It has been shown that representation quality is closely linked to generative performance (Yu et al., 2024). One line of work improves generation by regularizing model representations through alignment with strong pretrained networks. For example, Yu et al. (2024) aligns diffusion features with self-supervised features from DINOv2 (Oquab et al., 2023), while Wang et al. (2025) demonstrates that leveraging vision–language models (VLMs)  (Wang et al., 2024) can yield further improvements. Although such approaches accelerate training and enhance generation quality, they remain constrained by the need for high-quality external representations, which are not readily available in non-visual domains. Additionally, they introduce computational overhead, as pretrained models must be inferred at each training step. Another group of work adopts self-supervised strategies that rely on EMA (Exponential Moving Average)  (Tarvainen & Valpola, 2017) models to guide the representations. Zheng et al. (2023) integrates a generative diffusion process with an auxiliary mask reconstruction task. Zhu et al. (2024); Jiang et al. (2025) align representations between teacher and student encoders in a joint embedding space. While being self-contained, such methods increase computational cost by requiring an additional forward pass through the EMA model at each training step. Also, their performance still lags behind methods that leverage external supervision. A recent work (Wang & He, 2025) proposes dispersing representations in the feature space, analogous to the repulsive force in contrastive learning. This approach introduces no additional training overhead. Similarly, we present a self-contained, overhead-free solution; however, we leverage semantically richer internal representations to guide and improve the learning of weaker ones.
@@ 3.1 Preliminaries
We adopt the generalized perspective of stochastic interpolants (Ma et al., 2024),which provides a unifying framework for both flow-based and diffusion-based models. Here is a brief overview, we refer to the Appendix Appendix G for more details.
@@ 3.2 Diffusion Models Intermediate Representations
We investigate the representations learned by a pre-trained SiT model (Ma et al., 2024) on ImageNet (Deng et al., 2009). Through linear probing on downstream tasks (classification, segmentation) and Centered Kernel Alignment (CKA) with DINOv2 (Oquab et al., 2023) features, we observe a clear hierarchy in representation quality across layers. As shown in Figure 3(c), deeper layers exhibit superior discriminative capabilities, consistent with established principles of hierarchical feature learning in deep networks (LeCun et al., 2015). This pattern of increasing semantic richness culminating in a peak before the final decoding blocks is a known characteristic of diffusion model representations (Mukhopadhyay et al., 2024).
We propose to internally regularize the network by aligning representations from its early layers with those from its own deeper, semantically richer layers. Beyond the immediate benefit of improving shallow layers, we hypothesize that this method may induce a recursive refinement process. The enhancement of early-layer features is expected to facilitate the learning of more robust deep-layer representations, which subsequently offer a more refined target for the internal alignment, potentially leading to a cascading improvement of the model’s feature space.
@@ 3.3 LayerSync
We propose a self-contained regularization approach, named LayerSync, designed to improve the generation quality and training dynamics of diffusion models. The core principle behind LayerSync is intra-model self-alignment, where the model is trained to guide itself. We use the context-rich deep layers as an “intrinsic guidance” to provide a direct signal to the earlier “weak” layers, thereby enhancing the model’s entire feature from within.
@@
where the hyperparameter $\lambda>0$ balances the standard denoising task with our internal representation alignment. Algorithm 1 in Appendix summarizes our proposed approach.
@@ 3.4 Layer selection
A key design consideration for LayerSync is the selection of the layers to align. We note that this is a shared characteristic with other representation guidance methods, in particular those that rely on external supervision. We provide a simple yet effective heuristic that yields robust performance without requiring hyperparameter tuning.
Our selection strategy is guided by two principles. First, drawing on established findings (Xiang et al., 2023) regarding the functional specialization of layers in generative transformers, we exclude the final $20\%$ of blocks from being chosen as the reference layer as those are primarily specialized for low-level decoding tasks making them suboptimal as guidance targets. Second, to ensure a meaningful semantic gap between the representations, we enforce a minimum distance between the aligned and reference layers (e.g., 8 blocks for SiT-XL and 3 for SiT-B). We validate this heuristic and the robustness of the method to different layer selection through experiments as summarized in Table 5.
@@ 4.1 Image Generation
Implementation details. We strictly follow the setup in SiT  (Ma et al., 2024). Specifically, we use ImageNet  (Deng et al., 2009) and follow ADM  (Dhariwal & Nichol, 2021) for data preprocessing. The processed image will have the resolution of $256\times 256$ and is then encoded into a compressed vector $z\in\mathbb{R}^{32\times 32\times 4}$ using the Stable Diffusion VAE  (Rombach et al., 2022). For model configurations, we use the B/2, L/2, and XL/2 architectures by Ma et al. (2024), which process inputs with a patch size of 2. More details about the architectures and the number of parameters are provided in the Appendix J. We use cosine similarity between the patches as the similarity metric. Additional experimental details, including hyperparameter settings and computing resources, are provided in Appendix E.
@@
Evaluation metrics. We report Frechet inception distance (FID; Heusel et al. (2017)), Inception Score  (Salimans et al., 2016), Precision, and Recall  (Kynkäänniemi et al., 2019) using 50,000 samples. We provide details of each metric in Appendix H.
@@
Baselines. We compare our results with Dispersive  (Wang & He, 2025), the only self-contained, zero-cost method that accelerates training. For the sake of completeness, we also compare our method with several recent diffusion-based generation methods. For pixel-based approches we compare with ADM  (Dhariwal & Nichol, 2021), VDM++  (Kingma & Gao, 2023), Simple diffusion  (Hoogeboom et al., 2023), CDM  (Ho et al., 2022). For latent-based approaches we comapre with LDM  (Rombach et al., 2022), U-ViT-H/2  (Bao et al., 2023), DiffiT  (Hatamizadeh et al., 2024), MDTv2-XL/2  (Gao et al., 2023), MaskDiT  (Zheng et al., 2023), SD-DiT  (Zhu et al., 2024), DiT  (Peebles & Xie, 2023), and SiT  (Ma et al., 2024). We also compare our approach with REPA  (Zhang et al., 2025) and REED  (Wang et al., 2025) which rely on external representations. A detailed description of each baseline is provided in Appendix I.
@@
Results. As shown in Table 2, our method consistently improves diffusion transformer training and is more effective than Wang & He (2025). Our method results in $8.75\times$ acceleration compared to SiT-XL baseline, reaching an FID of 8.29 after only 160 epoches, and in $4.7\times$ acceleration compared to the baseline trained with Dispersive Loss. In table 2 we compare LayerSync with recent state-of-the-art diffusion model approaches. In particular, on SiTXL/2, we reach FID 1.89 after 800 epoches setting a new state-of-the-art in pure self-supervised generation, decreasing the gap with methods like Yu et al. (2024) that rely on external representations. We also qualitatively compare the progression of generation results in Figure 2, where we use the same initial noise across different models.
@@
Table 1: FID comparisons of class-conditional generation on ImageNet $\times256$. No classifier-free guidance (CFG; Ho & Salimans (2021)) is used. The sampler used is the ODE-based Heun method, except for the last section, which uses the SDE-based Euler method following Ma et al. (2024)
@@
Table 2: System-level comparison on ImageNet $\times256$ with CFG. Models with additional CFG scheduling are marked with an asterisk (*).
@@ 4.2 Audio Generation
Implementations details. We use the MTG-Jamendo dataset (Bogdanov et al., 2019), a large-scale collection containing over 55,000 full-length songs. For training, we process the audio by creating random 10-second samples, which are sampled at a standard rate of 44.1 kHz. We condition using the metadata provided with the dataset by conditioning the generation on the genre and instrument labels associated with each samples. Our audio generation model is an adaptation of the Scalable Interpolant Transformer (SiT-XL) (Ma et al., 2024), consistent with the 28-layer architecture used in our vision experiments. The model is configured to operate on patchified latent representations with a patch size of one. These latents are obtained from the pre-trained Variational Autoencoder (VAE) of the Stable Audio Open model (Evans et al., 2025), which provides a compact representation of the raw audio waveforms. The model was trained on 64 GH200 GPUs with a global batch size of 1024. In our experiment we align layer 8 with 21 using cosine similarity between the patch-wise representations of these two layers.
Evaluation metrics. To quantitatively assess the quality and realism of the generated audio, we report the Fréchet Audio Distance (FAD)(Kilgour et al., 2019) with 10.000 samples using the widely-used CLAP embeddings (Zhao et al., 2023).
@@
4.3 Text-condition Human Motion Generation

### v2

@@ 1 Introduction
Denoising generative models, such as diffusion  (Ho et al., 2020; Song et al., 2020; Song and Ermon, 2019) and flow matching models (Lipman et al., 2022), have demonstrated remarkable success in modeling complex data distributions, achieving state-of-the-art performance across a range of generative tasks. However, this success comes at a significant computation cost. Thus, a new promising line of research has emerged to improve the training efficiency of these models by improving the models’ intermediate representations  (Yu et al., 2024; Wang et al., 2025; Wang and He, 2025). It has been shown that the quality of a diffusion model’s intermediate representations is intrinsically linked to its generative performance. As a result, explicitly guiding these representations accelerates training and improves generation quality  (Yu et al., 2024).
@@
A recent step in this direction is the Dispersive Loss  (Wang and He, 2025), a self-contained regularizer that encourages internal representations to spread out in the feature space, analogous to the repulsive force in contrastive learning  (Oord et al., 2018). Although this approach demonstrates the potential for internal regularization, a substantial performance gap remains compared to methods that leverage external representations. In this paper, we propose a self-contained method with a more directed learning signal to reduce this gap.
Our work is motivated by two key observations: First, while diffusion models learn powerful representations, their quality is highly heterogeneous across the model’s depth. As demonstrated by previous works (Mukhopadhyay et al., 2024; Kim et al., 2025; Stracke et al., 2025), certain intermediate layers capture more semantically rich and useful information than others. Second, when models incorporate knowledge through external guidance using DINOv2 for instance, regularizing early layers seemed more effective than regularizing the deeper ones  (Yu et al., 2024; Wang et al., 2025). Based on these two observations, a clear opportunity presents itself: can the model’s own strongest layers act as an intrinsic guidance to improve its weaker ones through self-alignment?
To this end, we propose LayerSync, a simple yet powerful regularization framework that aligns a model’s own intermediate layers. LayerSync is a parameter-free, plug-and-play solution that operates without any external models or data, making it a truly self-contained method. LayerSync introduces negligible computational overhead, yet its effectiveness is substantial. Our experiments show that LayerSync consistently outperforms prior self-contained methods across all tested configurations. For image generation, LayerSync accelerates training on ImageNet 256×256 (Deng et al., 2009) by more than 8.75×. This leads to a new state-of-the-art in purely self-supervised image generation on ImageNet, demonstrating the strength of our self-alignment objective and substantially narrowing the gap between self-contained approaches and those relying on external guidance. Furthermore, due to its self-contained nature, LayerSync seamlessly generalizes to other modalities. Our experiments show that for audio generation, LayerSync leads to $21\%$ improvement in FAD-10K on MTG-Jamendo (Bogdanov et al., 2019), to $7.7\%$ improvement in FID for human motion generation on HumanML3D and $54.7\%$ in FVD for video generation on CLEVRER  (Yi et al., 2019) (see Appendix A). To the best of our knowledge, it is the first time that a self-contained method proves to accelerate diffusion model training seamlessly across different domains.
Additionally, an analysis of the internal features confirms that LayerSync strengthens the model’s representations, leading to a $32.4\%$ improvement in classification and a $63.3\%$ improvement in semantic segmentation accuracy.
@@
  • We demonstrate the domain-agnostic versatility of LayerSync by successfully applying it to image, audio, human motion, and video generation.
@@ 2 Related Work
Representation learning with diffusion models. Denoising Generative Models, including both diffusion  (Ho et al., 2020; Song et al., 2020; Song and Ermon, 2019) and flow matching models (Lipman et al., 2022), trained as multi-level denoising autoencoders (Vincent et al., 2008), naturally give rise to discriminative representations. A line of work has specifically evaluated the quality of these representations, showing that diffusion features can be effectively used across a variety of tasks  (Mukhopadhyay et al., 2024; Kim et al., 2025; Stracke et al., 2025) and, in some cases, achieve performance comparable to self-supervised representation learning methods  (Stracke et al., 2025). However, the quality of the representations varies across model layers, with the final layers, just before the model begins decoding, consistently containing more semantically rich features  (Kim et al., 2025; Xiang et al., 2023), regardless of whether the architecture is a U-Net  (Ronneberger et al., 2015) or a Transformer  (Vaswani et al., 2017). Our work is directly built upon those insights. We demonstrate that the semantically rich representations in the intermediate layers can be leveraged as a guidance signal to enhance the quality of earlier-layer representations.
@@
Representation regularization for improving diffusion models. It has been shown that representation quality is closely linked to generative performance (Yu et al., 2024). One line of work improves generation by regularizing model representations through alignment with strong pretrained networks. For example, Yu et al. (2024) aligns diffusion features with self-supervised features from DINOv2 (Oquab et al., 2023), while Wang et al. (2025) demonstrates that leveraging vision–language models (VLMs)  (Wang et al., 2024) can yield further improvements. Although such approaches accelerate training and enhance generation quality, they remain constrained by the need for high-quality external representations, which are not readily available in non-visual domains. Additionally, they introduce computational overhead, as pretrained models must be inferred at each training step. Another group of work adopts self-supervised strategies that rely on EMA (Exponential Moving Average)  (Tarvainen and Valpola, 2017) models to guide the representations. Zheng et al. (2023) integrates a generative diffusion process with an auxiliary mask reconstruction task. Zhu et al. (2024); Jiang et al. (2025) align representations between teacher and student encoders in a joint embedding space. While being self-contained, such methods increase computational cost by requiring an additional forward pass through the EMA model at each training step. A recent work (Wang and He, 2025) proposes dispersing representations in the feature space, analogous to the repulsive force in contrastive learning. This approach introduces no additional training overhead. Similarly, we present a self-contained, overhead-free solution; however, we leverage semantically richer internal representations to guide and improve the learning of weaker ones.
@@ 3.1 Preliminaries
We adopt the generalized perspective of stochastic interpolants (Ma et al., 2024), which provides a unifying framework for both flow-based and diffusion-based models. Here is a brief overview; we refer to Appendix Appendix N for more details.
@@ 3.2 Diffusion Models Intermediate Representations
Representation Hierarchy. We investigate the representations learned by a pre-trained SiT model (Ma et al., 2024) on ImageNet (Deng et al., 2009). Through linear probing on downstream tasks (classification, segmentation) and Centered Kernel Alignment (CKA) with DINOv2 (Oquab et al., 2023) features, we observe a clear hierarchy in representation quality across layers. As shown in Figure 4, deeper layers exhibit superior discriminative capabilities, consistent with established principles of hierarchical feature learning in deep networks (LeCun et al., 2015). This pattern of increasing semantic richness culminating in a peak before the final decoding blocks is a known characteristic of diffusion model representations (Mukhopadhyay et al., 2024).
Internal Block Structure of Diffusion Transformers. Beyond the hierarchy of feature quality, we observe that the internal structure of Diffusion Transformers at convergence follows specific correlation patterns as shown in Figure 2 and discussed in (Raghu et al., 2021; An et al., 2025). These blocks are highly correlated and naturally segregate into three functional groups: (1) an initial group focusing on local features, (2) a middle group of highly correlated blocks capturing global features, and (3) a final group acting as a decoder to project back to the latent space.
Figure 2: Correlation between the blocks of SiT-XL/2 at Convergence.
@@ 3.3 LayerSync
We propose a self-contained regularization approach, named LayerSync, designed to improve the generation quality and training dynamics of diffusion models. The core principle behind LayerSync is intra-model self-alignment, where the model is trained to guide itself. Building on the hierarchy of representation quality across layers, we use the context-rich deep layers as an “intrinsic guidance” to provide a direct signal to the earlier “weak” layers, thereby enhancing the model’s entire feature from within. The extra alignment between the model’s intermediate layers aligns with the block structure of the diffusion transformer at convergence.
@@
where the hyperparameter $\lambda>0$ balances the standard denoising task with our internal representation alignment. Algorithm 1 in the Appendix summarizes our proposed approach.
While LayerSync internally regularizes the network by aligning representations from its early layers with those from its own deeper, semantically richer layers, beyond the immediate benefit of improving shallow layers, we hypothesize that this method may induce a virtuous cycle in the recursive refinement process. The enhancement of early-layer features is expected to facilitate the learning of more robust deep-layer representations, which subsequently offer a more refined target for internal alignment, potentially leading to a cascading improvement of the model’s feature space.
@@ 3.4 Layer selection
A key design consideration for LayerSync is the selection of the layers to align. As detailed in Section 3.2, the layers of Diffusion Transformers naturally converge toward a structure in three groups of blocks focusing on local features, global features, and decoding. While our experiments indicate that LayerSync yields performance gains across a wide range of configurations, we observe that respecting this structure leads to optimal acceleration (see Appendix D).
Based on these observations, we propose a selection strategy guided by three principles. First, drawing on established findings (Xiang et al., 2023) regarding the functional specialization of layers in generative transformers, we exclude the final $20\%$ of blocks from being chosen as the reference layer, as those are primarily specialized for low-level decoding tasks, making them suboptimal as guidance targets. Second, based on the findings from (Raghu et al., 2021), we exclude the very first layers, as prior work finds that early blocks focusing on local features improve performance and generalization (An et al., 2025). Third, to ensure a meaningful semantic gap between the representations, we enforce a minimum distance between the aligned and reference layers (e.g., 8 blocks for SiT-XL and SiT-L and 3 for SiT-B). We validate this heuristic and the robustness of the method to different layer selection through experiments as summarized in Tables 6, 11 and 12.
@@ 4.1 Image Generation
Implementation details. We strictly follow the setup in SiT  (Ma et al., 2024). Specifically, we use ImageNet  (Deng et al., 2009) and follow ADM  (Dhariwal and Nichol, 2021) for data preprocessing. The processed image will have the resolution of $256\times 256$ and is then encoded into a compressed vector $z\in\mathbb{R}^{32\times 32\times 4}$ using the Stable Diffusion VAE  (Rombach et al., 2022). For model configurations, we use the B/2, L/2, and XL/2 architectures by Ma et al. (2024), which process inputs with a patch size of 2. More details about the architectures and the number of parameters are provided in the Appendix Q. Additional experimental details, including hyperparameter settings and computing resources, are provided in Appendix L.
@@
Evaluation metrics. We report Fréchet inception distance (FID; Heusel et al. (2017)), Inception Score  (Salimans et al., 2016), Precision, and Recall  (Kynkäänniemi et al., 2019) using 50,000 samples. We provide details of each metric in Appendix O.
@@
Baselines. We compare our results with Dispersive  (Wang and He, 2025), the only self-contained, zero-cost method that accelerates training. For the sake of completeness, we also compare our method with several recent diffusion-based generation methods. For pixel-based approaches we compare with ADM  (Dhariwal and Nichol, 2021), VDM++  (Kingma and Gao, 2023), Simple diffusion  (Hoogeboom et al., 2023), CDM  (Ho et al., 2022). For latent-based approaches we compare with LDM  (Rombach et al., 2022), U-ViT-H/2  (Bao et al., 2023), DiffiT  (Hatamizadeh et al., 2024), MDTv2-XL/2  (Gao et al., 2023), MaskDiT  (Zheng et al., 2023), SD-DiT  (Zhu et al., 2024), DiT  (Peebles and Xie, 2023), and SiT  (Ma et al., 2024). We also compare our approach with REPA  (Zhang et al., 2025) and REED  (Wang et al., 2025), which rely on external representations. Additionally, we compare to two autoregressive methods, VAR (Tian et al., 2024) and D-JEPA (Chen et al., 2024). For both methods, we selected the models with a similar number of parameters to SiT-XL. Finally, we compare our results with the EMA-based method SRA  (Jiang et al., 2025). A detailed description of each baseline is provided in Appendix P.
@@
Results. As shown in Table 2, our method consistently improves diffusion transformer training and is more effective than Wang and He (2025). Our method results in $8.75\times$ acceleration compared to SiT-XL baseline, reaching an FID of 8.29 after only 160 epochs, and in $4.7\times$ acceleration compared to the baseline trained with Dispersive Loss. In Table 2 we compare LayerSync with recent state-of-the-art diffusion model approaches. In particular, on SiT-XL/2, we reach FID 1.89 after 800 epochs, setting a new state-of-the-art in pure self-supervised generation, decreasing the gap with methods like Yu et al. (2024) that rely on external representations. We also qualitatively compare the progression of generation results in Figure 3, where we use the same initial noise across different models. Additional comparison metrics with SRA are provided in Table 15.
@@
Table 1: FID comparisons of class-conditional generation on ImageNet $\times256$. No classifier-free guidance (CFG; Ho and Salimans (2022)) is used. The sampler used is the ODE-based Heun method, except for the last section, which uses the SDE-based Euler method following Ma et al. (2024)
@@
Table 2: System-level comparison on ImageNet $\times256$ with CFG. Models with additional CFG scheduling Kynkäänniemi et al. (2024) are marked with an asterisk (*).
@@ 4.2 Audio Generation
Implementation details. We use the MTG-Jamendo dataset (Bogdanov et al., 2019), a large-scale collection containing over 55,000 full-length songs. For training, we process the audio by creating random 10-second samples, which are sampled at a standard rate of 44.1 kHz. We condition using the metadata provided with the dataset by conditioning the generation on the genre and instrument labels associated with each sample. Our audio generation model is an adaptation of the Scalable Interpolant Transformer (SiT-XL) (Ma et al., 2024), consistent with the 28-layer architecture used in our vision experiments. The model is configured to operate on patchified latent representations with a patch size of one. These latents are obtained from the pre-trained Variational Autoencoder (VAE) of the Stable Audio Open model (Evans et al., 2025), which provides a compact representation of the raw audio waveforms. The model was trained on 64 GH200 GPUs with a global batch size of 1024. In our experiment, we align layer 8 with 21 using cosine similarity between the patch-wise representations of these two layers.
Evaluation metrics. To quantitatively assess the quality and realism of the generated audio, we report the Fréchet Audio Distance (FAD)(Kilgour et al., 2018) with 10,000 samples using the widely-used CLAP embeddings (Zhao et al., 2023).
@@
4.3 Text-conditional Human Motion Generation